OpenStudy (anonymous):

The sum to infinity of G.P with common ratio r is 4 . The sum of the infinite G.P obtained by squaring the terms of this G.P is 16/3 then the first term and common ratio of given G.P is?

OpenStudy (dumbcow):

formula for sum of infinite GP is: \[\frac{a}{1-r}\] where a is first term and r<1 now set up 2 equations using given info \[\frac{a}{1-r} = 4\] \[\frac{a^{2}}{1-r} = \frac{16}{3}\] solve for a,r using substitution

OpenStudy (dumbcow):

no the common ratio must be less than 1 since the sum converges

OpenStudy (anonymous):

The answer should be a=2 r =1/2

OpenStudy (anonymous):

yup @dumbcow is correct

OpenStudy (dumbcow):

that answer doesn't work for 2nd sequence ?

hartnn (hartnn):

what @dumbcow ,said is correct,except a^2/(1-r^2)=16/3 did u get that??

OpenStudy (dumbcow):

oh yeah...oops the ratio is squared sorry

hartnn (hartnn):

u need explanation??i m getting r as 0.5,a=2......

OpenStudy (anonymous):

Yup

hartnn (hartnn):

a=4(r-1) \[a^2= 16/3*(r^2-1)\] squaring 1st equation and put a^2 in 2nd \[16(r-1)^2=16/3*(r^2-1)\] simplifying we get r=1/2

OpenStudy (anonymous):

but it should be (1-r)^2

hartnn (hartnn):

still u will get same result,i will be cancelling r-1 on both sides,u will be cancelling 1-r on both sides

OpenStudy (anonymous):

i mean the second eq

OpenStudy (anonymous):

The sum of the infinite G.P obtained by squaring the terms of this G.P is 16/3

OpenStudy (anonymous):

@hartnn

OpenStudy (anonymous):

a^2/(1-r)^2

hartnn (hartnn):

have u tried the same thing with (1-r)???i will write again using (1-r),but both things are essentially same..... a=4(1-r) a2=16/3∗(1-r2) squaring 1st equation and put a^2 in 2nd 16(1-r)2=16/3∗(1-r2) simplifying we get r=1/2

OpenStudy (anonymous):

it should be a^2=16/3∗(1-r)^2

OpenStudy (anonymous):

i think

hartnn (hartnn):

and no it should not be (1-r)^2 it must be (1-r^2) because the common ratio now is r^2 instead of r

OpenStudy (anonymous):

ok.....thxxxx

hartnn (hartnn):

1st term in new series is a^2,and common ratio is r^2 and formula is (1st term)/(1-common ratio)